# Systematic Classification of History-Dependent Invariants in the Lorenz System: Regularization Classes and Dynamical Signatures

### v1

@@ VIII.2 Physical Interpretation of the Conservation Laws
The history-dependent invariants occupy a distinctive position in the taxonomy of conservation laws.
Relation to classical conservation laws. In Hamiltonian mechanics, conservation laws arise from continuous symmetries via Noether’s theorem: energy from time-translation invariance, momentum from spatial translation, angular momentum from rotational invariance. These laws constrain motion to hypersurfaces in phase space and enable reduction of effective dimensionality.
The invariants $K_{n}=P_{n}+v_{n}$ share the formal property of constancy along trajectories but differ in origin and interpretation. They do not arise from a variational principle or symmetry of an action functional. Rather, they emerge from the algebraic structure of the Lorenz flow when augmented with auxiliary variables. Conservation is enforced by construction: $\dot{v}_{n}=-\dot{P}_{n}$ by definition.
@@
Non-triviality of the construction. A natural question is whether this conservation is merely tautological. The existence of the null class (Theorem 1) provides a negative answer. If conservation could be achieved for arbitrary polynomials $P(x,y,z)$ simply by defining $v=-\int\dot{P}\,dt$, then all 24 permutations would yield valid invariants. The failure of six permutations demonstrates that Lorenz structure genuinely constrains which polynomial forms are compatible with the orthogonality ansatz.
The null class failure reflects Schwarz integrability conditions imposing algebraic relations among partial derivatives that are violated when the auxiliary variable occupies the leading position. This obstruction is specific to the Lorenz vector field and would differ for other polynomial systems.
@@
First, the classification of compatible polynomial structures encodes information specific to the Lorenz dynamics. Eighteen valid invariants organized into three regularization classes reflect algebraic compatibility between the Lorenz flow and the orthogonality ansatz. Different classes correspond to different nullcline structures, connecting abstract algebraic classification to geometric organization of the flow.
Second, the distinct dynamical signatures exhibited by different classes provide complementary probes of attractor geometry. Statistical analysis demonstrates that Class III invariants function as geometric markers of topological transitions (lobe switches), while Class I invariants track continuous dynamics. This class-dependent sensitivity makes the classification a diagnostic tool for characterizing chaotic trajectories.
Operational interpretation. The invariants are most properly understood as constraints relating instantaneous configuration to accumulated trajectory history. While the form of conservation is guaranteed by $\dot{v}_{n}=-\dot{P}_{n}$, the existence of compatible polynomial structures $P_{n}$ reflects specific algebraic properties of the Lorenz equations.
VIII.3 Regularized Variables: Physical Interpretation
The central insight is that each invariant $K_{n}$ requires its own regularized variable $v_{n}(t)$, even when invariants belong to the same regularization class. Different conservation laws impose different demands on how trajectory history must be accumulated to maintain conservation.
The regularized variable $v_{n}(t)$ can be interpreted as the variation required to conserve $K_{n}$ along a trajectory. The evolution equation $\dot{v}_{n}=Q_{n}(x,y,z)$ specifies how this variation accumulates as the system evolves. The algebraic structure ensures that $v_{n}(t)=c_{n}-P_{n}$ fluctuates within a bounded range determined by attractor geometry. This variation is kinematic in nature, reflecting geometric accumulation of history required to satisfy the constraint $K_{n}=\mathrm{const}$. Unlike thermodynamic quantities (work, entropy production), $v_{n}$ carries no energetic content and does not satisfy a fluctuation-dissipation relation; it functions as a deterministic history accumulator, as discussed in Sec. II.2.
Statistical analysis (Sec. VII) quantifies the fluctuation patterns. Class III invariants incur large-amplitude spikes in $Q_{\mathrm{III}}$ at lobe-switching events, while Class I invariants accumulate more smoothly. This distinction reflects genuine differences in how invariants probe dynamics: Class III invariants are geometrically sensitive to topological transitions, while Class I invariants track more continuous aspects of the flow.
VIII.4 Relation to the Original Invariant
The invariant $K$ presented in Ref. [1] was derived using an orthogonality ansatz of the form $\nabla C=(-\dot{y},-\dot{u},\dot{z},-\dot{x})$, which assigns independent signs $\epsilon_{i}\in\{+1,-1\}$ to each permuted flow component. This represents one element of a larger family: allowing each of the four components to carry an independent sign yields $24\times 2^{4}/2=192$ distinct ansätze (the factor of $1/2$ accounts for the global gauge symmetry $K\to-K$). The systematic exploration presented here restricts attention to the 24 ansätze with uniform global sign, i.e., $\nabla C=\pm\pi(\mathbf{f})$.
Explicit mapping. The ansatz of Ref. [1] corresponds to the permutation $(y,u,z,x)\to(2,4,3,1)$ with sign pattern $(\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4})=(-1,-1,+1,-1)$. However, the regularization polynomial determines class membership, not permutation indices. The original invariant $\mathcal{K}$ has the form (Eq. 8 of Ref. [1]):
$\mathcal{K}=xy(z+1)+\tfrac{1}{2}x^{2}(z-\rho)-\tfrac{1}{2}\beta z^{2}+\sigma v,\quad v=(x-y)u.$ (78)
The regularized variable v=(xy)uv=(x-y)u absorbs the singularity at $x=y$, identifying the regularization polynomial as $(x-y)=-p_{\mathrm{I}}$. Therefore, the original invariant $\mathcal{K}$ belongs to Class I, sharing singularity structure with $K_{1}$$K_{6}$. Despite the non-uniform sign pattern, $\mathcal{K}$ exhibits the same geometric sensitivity as Class I invariants: its evolution function $Q=\dot{v}$ varies smoothly across the attractor without the large spikes characteristic of Class III. This confirms that the regularization class structure, determined by which nullcline removes the singularity, is more fundamental than the specific permutation or sign convention.
Both families share the same regularization structure: the singularity-removing polynomials $p_{\mathrm{I}}=y-x$, $p_{\mathrm{II}}=y+x(z-\rho)$, and $p_{\mathrm{III}}=xy-\beta z$ appear in both constructions. This shared structure suggests that the class organization is a robust feature of the Lorenz system.
VIII.5 Operational Interpretation and Finite Memory Horizons
The history-dependent nature of the invariants raises a practical question: what is the operational meaning of a conserved quantity that depends on the infinite past of a trajectory?
The regularized variable $v_{n}(t)$, defined through $v_{n}(t)=v_{n}(0)+\int_{0}^{t}Q_{n}\,ds$, accumulates trajectory history from $t=0$. In the deterministic setting, this accumulation is exact and $K_{n}$ remains conserved. However, in any physical realization, small perturbations (numerical errors, thermal fluctuations, external noise) cause trajectory deviations. Since $v_{n}(t)$ integrates over all time, even small persistent perturbations can lead to secular drift.
A natural resolution interprets the invariants as cycle-averaged quantities. Consider a Poincaré section $\Sigma$ defined by a transverse surface, for instance $z=\rho-1$. Each time the trajectory crosses $\Sigma$, we reset $v_{n}\to 0$. The value accumulated during one return defines production of history for that cycle:
$\Delta v_{n}=\oint_{\text{cycle}}Q_{n}(x,y,z)\,dt.$ (79)
This cycle-averaged interpretation admits a rigorous foundation. For UPOs, $\Delta v_{n}$ takes a fixed value characteristic of the orbit, providing the same fingerprint discussed earlier. For chaotic trajectories, $\{\Delta v_{n}\}$ forms a discrete-time observable whose statistical properties are governed by the ergodic structure of the attractor.
VIII.6 Robustness Under Stochastic Perturbations
The deterministic conservation laws $dK_{n}/dt=0$ hold exactly when the system evolves according to the Lorenz equations (1)–(3). In the presence of additive noise, a more realistic model of physical systems, the evolution becomes
$dx$ $=\sigma(y-x)\,dt+\epsilon\,dW_{x},$ (80)
$dy$ $=[x(\rho-z)-y]\,dt+\epsilon\,dW_{y},$ (81)
$dz$ $=(xy-\beta z)\,dt+\epsilon\,dW_{z},$ (82)
where $dW_{i}$ are independent Wiener increments and $\epsilon$ measures noise strength.
Under such perturbations, the exact cancellation that ensures $dK_{n}/dt=0$ fails. The invariant $K_{n}$ becomes a stochastic process:
$dK_{n}=\epsilon\sum_{i}\frac{\partial K_{n}}{\partial x_{i}}dW_{i},$ (83)
a random walk with variance growing as $\epsilon^{2}t$.
The characteristic time over which invariants remain approximately conserved scales as:
$\tau_{\text{useful}}\sim\frac{1}{\epsilon^{2}\langle\|\nabla_{xyz}K_{n}\|^{2}\rangle_{\mathrm{SRB}}},$ (84)
where $\langle\cdot\rangle_{\mathrm{SRB}}$ denotes the average over the SRB measure. The quadratic sensitivity to noise distinguishes history-dependent invariants from topological invariants (such as knot types of periodic orbits), which are robust under arbitrary continuous deformations.
Diagnostic utility of structural redundancy. The existence of 18 algebraically distinct invariants with functional rank 4 provides internal consistency checks. For any pair of invariants within the same class, the identity $v_{n}(t)-v_{m}(t)=P_{m}(x,y,z)-P_{n}(x,y,z)$ must hold exactly in the deterministic limit, and to within Brownian diffusion under stochastic perturbations. A sudden violation of this identity, beyond the expected $O(\epsilon\sqrt{t})$ drift, signals numerical integration failure rather than physical dynamics. This is particularly valuable near nullcline crossings where stiffness is high. Furthermore, comparing invariants from different classes provides complementary information: Class III invariants signal topological transitions (lobe switches), while Class I invariants track continuous dynamics. Under noise, a spike in $Q_{\mathrm{III}}$ accompanied by smooth $Q_{\mathrm{I}}$ behavior confirms a genuine separatrix crossing, whereas uncorrelated fluctuations in both suggest stochastic perturbation. The structural redundancy thus functions as a filter bank for distinguishing topology from noise.
VIII.7 Numerical Stiffness and Practical Integration
While the regularization procedure (Sec. III.4) ensures that all invariants are theoretically well-defined throughout phase space, practical numerical integration of Class III evolution functions presents challenges.
The evolution equation $\dot{v}_{n}=Q_{n}(x,y,z)$ requires evaluating the polynomial $Q_{n}$ along the trajectory. For Class III invariants, statistical analysis reveals that $|Q_{\mathrm{III}}|$ reaches values of order $10^{4}$$10^{5}$ during lobe-switching events, while remaining of order $10^{0}$$10^{1}$ between transitions. This disparity constitutes numerical stiffness.
The stiffness ratio can be estimated as:
$\mathcal{S}\equiv\frac{|Q_{n}|_{\text{max}}}{|Q_{n}|_{\text{typical}}}\sim\frac{T}{\tau_{\text{cross}}}\approx 10^{3}\text{--}10^{4},$ (85)
where $T\approx 1.5$ is the typical orbital period and $\tau_{\text{cross}}\approx 10^{-3}$$10^{-4}$ is the separatrix crossing time.
Three practical implications follow:
@@ VIII.7 Numerical Stiffness and Practical Integration
    Adaptive step control: The StiffnessSwitching method employed in validation automatically detects stiff intervals and adjusts the integration algorithm. This is essential for maintaining accuracy during Class III spikes.
@@
    Class-dependent accuracy: Class I and Class II invariants, lacking the large spikes of Class III, can be integrated with coarser timesteps. For applications requiring only Class I fingerprints, computational costs are substantially reduced.
@@
    Spike-aware sampling: When computing statistics of $Q_{n}(t)$, fixed-interval sampling may undersample spike events. The heavy tails in Fig. 2(b) are accurately captured only with sufficiently fine temporal resolution ($\Delta t\ll\tau_{\text{cross}}$).
The distinction between analytical smoothness and numerical stiffness is important: invariants are mathematically well-defined everywhere, but practical computation requires integration schemes capable of handling large derivative ratios. This is a standard challenge in stiff ODE systems and does not reflect any deficiency in the invariant construction.
VIII.8 Forward and Inverse Problems
A complete assessment of the invariants’ utility requires distinguishing two complementary perspectives.
The Forward Problem. In the forward or initial value problem, one specifies initial conditions $(x_{0},y_{0},z_{0})$ and asks: where does the trajectory go? Since $v_{n}(0)=0$ for all $n$, the invariant values are determined instantaneously:
$c_{n}=K_{n}(0)=P_{n}(x_{0},y_{0},z_{0}).$ (86)
The Triad fingerprint $(c_{1},c_{2},c_{3})$ is determined algebraically from initial conditions without integration. The fingerprint is an a priori property of the trajectory.
The Inverse Problem. In the inverse or classification problem, one specifies a Triad of invariant values $(c_{1},c_{2},c_{3})$ and asks: what initial conditions are compatible? The three polynomial equations $P_{n}(x,y,z)=c_{n}$ generically determine a discrete set of candidate initial states. This provides a classification tool: periodic orbits can be catalogued by fingerprints, and chaotic trajectories can be organized into families based on invariant values.
VIII.9 From Algebra to Dynamics
The partition into three classes based on regularization polynomial $p_{n}$ was established algebraically. Statistical analysis demonstrates genuine dynamical consequences, elevating the classification from mathematical convenience to physically meaningful distinction.
The class structure reflects different response of $Q_{n}(t)$ to trajectory motion:
  • Class I (regularized by $p_{\mathrm{I}}=y-x$): The nullcline $y=x$ is crossed at various orbital phases, not specifically at lobe-switching events. This explains moderate kurtosis ($\kappa_{\mathrm{I}}\approx 8$).
  • Class III (regularized by $p_{\mathrm{III}}=xy-\beta z$): Polynomial parts contain terms proportional to $xy$, which changes sign at separatrix crossings. Class III invariants are sensitive to topological transitions. Elevated kurtosis ($\kappa_{\mathrm{III}}\approx 15$) and strong negative skewness ($S_{\mathrm{III}}\approx-2.7$) reflect the concentration of large negative spikes at lobe switches, where transitions occur in the upper region of the attractor.
The spikes in $Q_{\mathrm{III}}(t)$ function as markers signaling each addition to the symbolic sequence $\ldots LLRLLRLR\ldots$. Class III invariants serve as geometric probes of topological transitions.
@@ IX Conclusion
This work has presented a systematic exploration of history-dependent dynamical invariants in the Lorenz system using permutation-based enumeration. The principal results are:
1. Enumeration: Of 24 possible permutation ansätze with uniform global sign, exactly 18 yield valid invariants ($K_{1}$$K_{18}$). The failure of 6 permutations demonstrates that Lorenz structure constrains which polynomial forms can be conserved; the construction is algebraically constrained, not arbitrary.
2. Classification: The invariants organize into three classes (I, II, III) determined by the regularization polynomial, plus a null class (4abc) whose failure reflects Schwarz integrability conditions of the augmented system.
3. Regularized variables: Each invariant $K_{n}$ requires its own regularized variable $v_{n}(t)$ satisfying $\dot{v}_{n}=Q_{n}(x,y,z)$. Regularized variables remain bounded, as established by $v_{n}=c_{n}-P_{n}$ and compactness of the attractor. Original auxiliary variables $u_{n}=v_{n}/p_{n}$ diverge at nullcline crossings, representing coordinate singularities that do not affect invariant conservation.
4. Independence of history constraints: Any selection of three invariants from different regularization classes provides independent constraints on trajectory histories. The Triad fingerprint $(c_{1},c_{2},c_{3})$ characterizes each trajectory up to the discrete $\mathbb{Z}_{2}$ symmetry. Parity-breaking Triads (containing odd invariants) can distinguish symmetric trajectories.
5. Immediate determination: The Triad fingerprint is determined instantaneously from initial conditions via $c_{n}=P_{n}(x_{0},y_{0},z_{0})$, since $v_{n}(0)=0$. The constraints are a priori properties of trajectories.
6. Class dynamics: The three regularization classes exhibit distinct dynamical signatures in the chaotic regime. Class III displays higher intermittency (kurtosis $\kappa_{\mathrm{III}}\approx 15$ versus $\kappa_{\mathrm{I}}\approx 8$; Fig. 2b) and stronger negative skewness ($S_{\mathrm{III}}\approx-2.7$ versus $S_{\mathrm{I}}\approx-1.6$), correlating with lobe-switching events in the upper attractor region. This divergence is absent pre-chaos (Fig. 2a), confirming it arises from the interaction between polynomial structure and chaotic dynamics. Class III evolution functions $Q_{\mathrm{III}}(t)$ exhibit negative spikes (Fig. 2c) that encode the symbolic dynamics of the chaotic flow.
7. Physical interpretation: The regularized variable $v_{n}(t)$ represents the kinematic variation required to conserve $K_{n}$ along a trajectory. Different invariants have different variation functions, and the statistical properties of these variations distinguish continuous-tracking (Class I) from topological-transition-sensitive (Class III) invariants.
8. Utility: The invariants provide: (i) independent monitors of numerical integration accuracy, (ii) unique fingerprints for classifying unstable periodic orbits, (iii) trajectory-discriminating quantities that encode information about the initial state not recoverable from the current position alone, (iv) complementary probes for characterizing attractor structure, with Class III invariants serving as geometric probes of topological transitions while Class I invariants track continuous dynamics, and (v) a conceptual framework relating chaos to incompatibility of different measures of trajectory history.
9. Relation to original invariant: The invariant $\mathcal{K}$ of Ref. [1] employs a non-uniform sign ansatz belonging to the complementary 192-element family. The regularized variable v=(xy)uv=(x-y)u identifies the regularization polynomial as $(x-y)=-p_{\mathrm{I}}$, placing $\mathcal{K}$ in Class I. This confirms that class structure is determined by the nullcline that removes the singularity, not by permutation indices or sign pattern.
Open questions: Several problems emerge from this analysis: (i) What is the complete structure of the 192-element family with independent component signs? (ii) Does this framework extend to other dissipative chaotic systems, and what determines invariant family structure? (iii) Can the bounded fluctuations of regularized variable differences $v_{n}(t)-v_{m}(t)$ be exploited as diagnostics of attractor geometry? (iv) What is the explicit form of the probability density function of cycle-averaged $\Delta v_{n}$, and how does its structure relate quantitatively to attractor symbolic dynamics? (v) In stochastic settings, can dynamic selection among invariants with locally minimal $\|\nabla P_{n}\|$ extend the useful prediction horizon beyond the estimate (84)?
The systematic framework established here, comprising permutation-based enumeration with regularization analysis and statistical characterization, is applicable to other dissipative systems and may reveal similar hidden structure in chaotic dynamics.

### v2

@@ VIII.2 Physical Interpretation of the Conservation Laws
The history-dependent invariants occupy a distinctive position in the taxonomy of conservation laws. They do not arise from a variational principle or symmetry of an action functional; rather, they emerge from the algebraic structure of the Lorenz flow when augmented with auxiliary variables. Conservation is enforced by construction: $\dot{v}_{n}=-\dot{P}_{n}$ by definition.
@@
Non-triviality. The existence of the null class demonstrates that conservation cannot be achieved for arbitrary polynomials. The failure of six permutations to yield valid invariants proves that the Lorenz structure genuinely constrains which polynomial forms are compatible with the orthogonality ansatz.
@@
    The classification of compatible polynomial structures encodes information specific to the Lorenz dynamics. Eighteen valid invariants organized into three regularization classes reflect algebraic compatibility between the Lorenz flow and the orthogonality ansatz.
@@
    The distinct dynamical signatures exhibited by different classes provide complementary probes of attractor geometry. Class III invariants function as geometric markers of topological transitions, while Class I invariants track continuous dynamics.
VIII.3 Connection to Template Theory
The connection between Class III sensitivity and lobe-switching events resonates with the template theory of chaotic attractors [6]. The Lorenz attractor possesses a branched manifold structure that captures the essential stretching and folding operations; periodic orbits can be classified by their knot type, determined by the sequence of lobe visits.
The evolution function $Q_{\mathrm{III}}(t)$ is functionally equivalent to a continuous Poincaré section detector: each spike marks a crossing of the template’s branch line. The integrated area under each spike corresponds to the “topological contribution” of that symbol transition, providing a continuous encoding of the discrete symbolic dynamics.
The scaling law $\Delta t=t_{\min}+C\mathcal{A}^{-n}$ with $n\propto\beta$ establishes a quantitative link between the algebraic structure of the invariants and the eigenvalue structure of the flow. The stable eigenvalue $\beta$ governs relaxation along the $z$-direction, and the Class III regularization polynomial $p_{\mathrm{III}}=xy-\beta z$ explicitly couples to this mode. The linear relationship $n=m\beta+b$ with $m\approx 0.2$ indicates that the detector integrates information over multiple relaxation cycles.
VIII.4 Robustness and Practical Utility
The invariants provide several practical utilities:
  1. 1.
    Numerical accuracy monitors: Conservation violations signal integration errors.
  2. 2.
    Periodic orbit fingerprints: The Triad $(c_{1},c_{2},c_{3})$ provides intrinsic orbit labels.
@@ VIII.4 Robustness and Practical Utility
    Trajectory classification: Different initial conditions yield different fingerprints.
  4. 4.
    Topological precursor detection: The differential signal $\Delta S$ provides advance warning of separatrix crossings.
The structural redundancy of 18 invariants provides internal consistency checks. For pairs within the same class, $v_{n}-v_{m}=P_{m}-P_{n}$ must hold exactly. Comparing invariants from different classes enables filtering topology from noise: a spike in $Q_{\mathrm{III}}$ accompanied by smooth $Q_{\mathrm{I}}$ behavior confirms a genuine separatrix crossing, whereas uncorrelated fluctuations suggest stochastic perturbation.
VIII.5 Experimental Realization and Long-Time Stability
While the differential detector (Sec. VII.4) is optimal for short-term prediction of imminent separatrix crossings, practical implementation in physical experiments or long numerical simulations requires addressing two fundamental challenges that arise from the history-dependent nature of the invariants.
The secular drift problem. Since $v_{n}(t)$ integrates the trajectory history via $v_{n}=\int_{0}^{t}Q_{n}\,ds$, any small but persistent perturbation (thermal noise, roundoff error, measurement uncertainty) accumulates over time. In an open-loop implementation, this produces a secular drift: the invariant ceases to be constant and executes a random walk with variance growing as $\sigma^{2}\sim t$, rendering long-term conservation unusable.
The unknown initial condition problem. Computing the conserved value $c_{n}=P_{n}(x_{0},y_{0},z_{0})$ requires knowledge of the exact initial state. In chaotic experimental systems, information about the initial condition is lost exponentially fast. An observer who begins measurement at time $t>0$ cannot determine the invariant value without access to the prior history.
Both problems are resolved by the operational reset protocol. Rather than treating $v_{n}$ as an open-ended integral accumulating from $t=0$ to the current time $t$, one selects a Poincaré section $\Sigma$ transverse to the flow (for example, the plane $x=0$ with $\dot{x}>0$). Each time the trajectory crosses $\Sigma$, the accumulated variable is reset: $v_{n}\to 0$. The physically meaningful quantity becomes the discrete sequence of cycle-integrated values
$\Delta v_{n}^{(k)}=\int_{t_{k}}^{t_{k+1}}Q_{n}(x(s),y(s),z(s))\,ds,$ (57)
where $t_{k}$ denotes the $k$-th crossing of $\Sigma$.
This cyclic formulation transforms the open-loop integral (unbounded, drift-prone) into a closed-loop measurement (bounded, self-correcting). For unstable periodic orbits, the value $\Delta v_{n}^{(\mathrm{UPO})}$ after one complete period is a fixed topological-dynamical invariant, providing an unambiguous orbit fingerprint independent of when measurement begins. The statistical distribution of $\{\Delta v_{n}^{(k)}\}$ across many cycles encodes intrinsic properties of the attractor’s ergodic structure, connecting the continuous dynamics to discrete symbolic analysis.
@@ IX Conclusion
This work has presented a systematic exploration of history-dependent dynamical invariants in the Lorenz system and demonstrated their utility as deterministic probes of topological transitions. The investigation began with a methodical question: given that the orthogonality construction admits multiple polynomial forms, how many independent invariants exist, and do they possess distinct dynamical signatures? The answer reveals an unexpectedly rich algebraic structure with direct physical consequences.
The permutation-based enumeration of the augmented phase space yields a complete classification. Of the 24 possible ansätze, exactly 18 produce valid invariants organized into three regularization classes, while the remaining 6 fail due to Schwarz integrability constraints. This partition is not arbitrary; the first coordinate in each permutation determines the regularization polynomial, which in turn controls the geometric coupling between the invariant and the Lorenz flow. The null class, where the auxiliary variable occupies the leading position, embodies a fundamental asymmetry: the memory kernel must depend on the resolved variables, not vice versa.
The central discovery concerns the operational distinction between regularization classes. Statistical analysis of $N=4\times 10^{6}$ trajectory samples demonstrates that Class III invariants, whose polynomial parts contain the bilinear product $xy$, function as geometric probes of lobe-switching events. In the chaotic regime, Class III exhibits elevated kurtosis ($\kappa_{\mathrm{III}}\approx 15.8$ versus $\kappa_{\mathrm{I}}\approx 8.2$) with heavy tails synchronized to separatrix crossings, while both classes remain statistically similar in the pre-chaotic regime. This divergence confirms that the signatures arise specifically from the interaction between polynomial structure and chaotic dynamics, rather than from any intrinsic bias in the formulation.
The complementarity between classes enables construction of a differential detector $\Delta S=Q_{13}-Q_{5}$ that provides advance warning of topological transitions. The prediction horizon follows a shifted power law $\Delta t=t_{\min}+C\mathcal{A}^{-n}$, where parametric analysis across $\beta\in[2.0,3.2]$ reveals that the exponent scales linearly with the stable eigenvalue: $n=m\beta+b$ with $m=0.196\pm 0.02$ and $b\approx 0.5$, with $R^{2}=0.94$. This quantitative link between algebraic classification and predictive capability establishes that the Class III regularization polynomial $p_{\mathrm{III}}=xy-\beta z$ explicitly couples to the relaxation mode governing separatrix approach.
High-precision numerical validation using 80-digit arithmetic confirms conservation to $O(10^{-36})$, demonstrating that the invariants are exact within the limits of numerical representation. For practical implementation in physical experiments or long numerical simulations, the operational reset protocol transforms the history-dependent integrals into cycle-bounded measurements anchored to a Poincaré section. This cyclic formulation eliminates secular drift and enables robust characterization even when initial conditions are unknown.
The framework established here opens several avenues for future investigation. Whether the linear scaling $n\propto\beta$ holds for other sensor pairs from different classes remains to be tested. The complete structure of the 192-element family obtained by allowing independent component signs has not been explored. Most significantly, the question of whether analogous history-dependent invariants exist in other dissipative chaotic systems suggests that the Lorenz construction may represent a specific instance of a more general phenomenon. The systematic methodology developed here, comprising permutation-based enumeration with statistical characterization and topological precursor analysis, provides a template for such investigations.